Circular No. 8102

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003co Further to IAUC 8101, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of an apparent supernova (mag 18.6) on unfiltered Palomar NEAT images taken on Mar. 22.14, 22.16, and 22.18 UT. The new object is located at α = 10^{h}49^{m}39^{s}.91, δ = +29^{o}21′44′′.6 (equinox 2000.0), which is 7′′.4 west and 7′′.9 south of the center of the apparent host galaxy. Additional magnitudes of SN 2003co from NEAT images: 2002 Mar. 31.32, [20.9; 2003 Feb. 21.46, [20.9; Mar. 30.14, 18.1.

XTE J1550–564 P. Woudt, University of Cape Town (UCT); and P. Charles and I. C.

Shih, University of Southampton, report that, following the INTEGRAL detection of renewed hard x-ray activity in the blackhole transient XTE J1550–564 (cf. IAUC 8100), they obtained high-speed photometry of the optical counterpart (Jain et al. 2001, Ap.J. 554, L181) with the South African Astronomical Observatory 1.9-m telescope (+ UCT CCD) during 25 min on Mar. 27. The object appeared at V ∼ 18 (close to its maximum during the 2000 outburst) and it displays significant variability on timescales of 1–2 min, including a flare of ∼ 0.5 mag lasting < 20 s. Further infrared and radio observations of this microquasar are encouraged.

GRB 030329 J. E. McGaha, Tucson, AZ, reports that CCD images taken with a 0.30-

m f/10 Schmidt-Cassegrain reflector show an object at α = 10^{h}44^{m}50^{s}.01, δ = +21^{o}31′17′′.8 (equinox 2000.0) that faded from red mag 16.2 on Mar. 31.201 UT to 16.4 on Mar. 31.289. Nothing is visible at this location on a Digitized Sky Survey plate taken in 1996 (limiting mag ∼ 20.5).

SUPERNOVA 2003cn IN IC 849 M. Hamuy, Carnegie Observatories, reports that a spectrum (range

380–730 nm) of SN 2003cn (cf. IAUC 8101), obtained on Mar. 31.31 UT with the Las Campanas Dupont 2.5-m telescope, shows that this is a type-II supernova characterized by Balmer lines exhibiting P-Cyg profiles, although the absorption component of Hα is very weak. The expansion velocity of 5900 km/s, derived from the minimum of Hβ (assuming the NED recession velocity of 5430 km/s for the host galaxy), and the weak He I 587.6-nm line suggest that SN 2003cn is ∼ 2–3 weeks old.

2003 March 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green